jojo's java hut mixed Kona coffee beans worth$6/pound with columbian beans with $8.80/pound. the resulting mixture was 16 pounds of a delightful breakfast blend worth $6.70/pound. how many pounds of each type of bean were used?
Answer by scott8148(6628)

k + c = 16

k(6) + c(8.8) = 16(6.7)